OpenStudy (anonymous):

what is the factor of 64x3-8 then solve it?

OpenStudy (zzr0ck3r):

(64x^3) -8 = 4^3*x^3-2^3 = (4x)^3 - 2^3 I think you have a formula for the rest?

OpenStudy (zzr0ck3r):

hint: a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (anonymous):

can you solve the given factor?

OpenStudy (zzr0ck3r):

I have showed you how you have (4x)^3-2^3 and a^3-b^3 = (a-b)(a^2+ab+b^2) so let a = 4x and b= 2 what do you have?

OpenStudy (anonymous):

can you give me on how to check my answer if its right?